# LogicMP: A Neuro-symbolic Approach for Encoding First-order Logic Constraints

### v3

@@ 3 Efficient Mean-field Iteration via LogicMP
$\phi_{f}(v_{g})$ $v_{\mathtt{C}(e_{1},e_{3})}$ useful
00 $1$
$g_{-\mathtt{C}(e_{1},e_{3})}$ $(0,0)$ $1=1$
@@ 3.2 Parallel Aggregation using Einstein Summation
$\check{\mathbf{Q}}^{[f,h]}_{r_{h}}(\mathbf{v}_{r_{h}})\leftarrow\bm{1}_{\mathbf{v}_{r_{h}}=\neg n_{h}}\mathtt{einsum}(\text{``}...,\mathcal{A}_{r_{j\neq h}}^{f},...\to\mathcal{A}_{r_{h}}^{f}\text{''},...,\mathbf{Q}_{r_{j\neq h}}(n_{j\neq h}),...)\,,$ (4)
@@ 5.2 Encoding FOLCs over Relational Graphs
MLN MCMC (Richardson & Domingos, 2006) .53 - - - - - - - - - - - - - - - - -
@@ 5.3 Encoding FOLCs over Text
[…]
@@ Appendix A Mean-field Iteration Equation of MLN Inference
$\begin{split}D_{KL}(Q(\mathbf{v})||p(\mathbf{v}|O))&=\mathbb{E}_{Q(\mathbf{v})}\log\frac{Q(\mathbf{v})}{p(\mathbf{v}|O)}\\ &=\mathbb{E}_{Q(\mathbf{v})}\log{Q(\mathbf{v})}-\mathbb{E}_{Q(\mathbf{v})}{\log p(\mathbf{v}|O)}\\ &=\mathbb{E}_{Q(\mathbf{v})}\log{Q(\mathbf{v})}-\mathbb{E}_{Q(\mathbf{v})}{\log p(\mathbf{v},O)}+\log p(O)\\ &=-(\mathbb{E}_{Q(\mathbf{v})}{\log p(\mathbf{v},O)}-\mathbb{E}_{Q(\mathbf{v})}\log{Q(\mathbf{v})})+\log p(O)\,,\\ \end{split}$ (7)
@@
$\begin{split}\mathbb{E}_{Q(\mathbf{v})}\log{p(\mathbf{v}|O)}&=\sum_{i,v_{i}}\phi_{u}(v_{i})Q_{i}(v_{i})+\sum_{f\in F}w_{f}\sum_{g\in G_{f}}\sum_{\textbf{v}_{g}}\phi_{f}(\textbf{v}_{g})\prod_{i\in g}Q_{i}(v_{i})-\log Z\,,\end{split}$ (9)
@@
$\begin{split}\mathcal{L}_{i}&=\sum_{v_{i}}Q_{i}(v_{i})\log Q_{i}(v_{i})\\ &-\sum_{v_{i}}Q_{i}(v_{i})[\phi_{u}(v_{i})+\sum_{f\in F}w_{f}\sum_{g\in G_{f}(i)}\sum_{\textbf{v}_{g_{-i}}}\phi_{f}(v_{i},\textbf{v}_{g_{-i}})\prod_{j\in g_{-i}}Q_{j}(v_{i})]\,,\\ \end{split}$ (11)
@@ Appendix B Terminology
Clause. A clause is a formula formed from a finite disjunction of literals. A clause is true whenever at least one of the literals that form it is true. Clauses are usually written as $(l_{1}\vee l_{2}...)$, where the symbols $l_{i}$ are literals, e.g., $\mathtt{S}(a)\wedge\mathtt{F}(a,b)\implies\mathtt{S}(b)$.
@@
Conjunctive Normal Form (CNF). A formula is in conjunctive normal form if it is a conjunction of one or more clauses, where a clause is a disjunction of literals. The CNF is usually written as $(l_{1,1}\vee l_{1,2}...)\wedge(l_{2,1}\vee l_{2,2}...)\wedge...$, where the symbols $l_{i,j}$ are literals, e.g., $(\mathtt{S}(a)\implies\mathtt{C}(a))\wedge(\mathtt{C}(a)\implies\mathtt{S}(a))$.
@@ Appendix C Proof of Theorem: Message of Clause Considers True Premise Only
$\begin{split}Q_{i}(v_{i})&=\frac{\exp(E_{i}(v_{i}))}{\sum_{v_{i}}\exp(E_{i}(v_{i}))}\,,\\ E_{i}(v_{i})&=\phi_{f}(v_{i},\mathbf{v}^{*}_{g^{*}_{-i}};\mathbf{n})\prod_{j\in{g^{*}_{-i}}}Q_{j}(v^{*}_{j})+\Delta_{i}(v_{i})\,,\\ \Delta_{i}(v_{i})&=\phi_{u}(v_{i})+\sum_{\mathbf{v}_{g^{*}_{-i}}{\neq\mathbf{v}^{*}_{g^{*}_{-i}}}}\phi_{f}(v_{i},\mathbf{v}_{g^{*}_{-i}};\mathbf{n})\prod_{j\in{g^{*}_{-i}}}Q_{j}(v_{j})\\ &+\sum_{f}w_{f}\sum_{g\in G_{f}(i)\backslash g^{*}}\sum_{\mathbf{v}_{g_{-i}}}\phi_{f}(v_{i},\mathbf{v}_{g_{-i}};\mathbf{n}^{f})\prod_{j\in{g_{-i}}}Q_{j}(v_{j}))\,.\\ \end{split}$ (17)
@@ Appendix D Proof of Theorem: Message of CNF = $\sum$ Message of Clause
$\begin{split}Q_{i}(v_{i})&=\frac{\exp(E_{i}(v_{i}))}{\sum_{v_{i}}\exp(E_{i}(v_{i}))}\,,\\ E_{i}(v_{i})&=\sum_{\mathbf{v}_{g^{*}_{-i}}}\phi_{{f}}(v_{i},\mathbf{v}_{g^{*}_{-i}})\prod_{j\in{g^{*}_{-i}}}Q_{j}(v_{j})+\Delta_{i}(v_{i})\,,\\ \Delta_{i}(v_{i})&=\phi_{u}(v_{i})+\sum_{f}w_{f}\sum_{g\in G_{f}(i)\backslash g^{*}}\sum_{\mathbf{v}_{g_{-i}}}\phi_{f}(v_{i},\mathbf{v}_{g_{-i}})\prod_{j\in{g^{*}_{-i}}}Q_{j}(v_{j}))\,.\\ \end{split}$ (19)
@@ Appendix G Einsum Examples
[…]
@@ Appendix I An Example to Illustrate the LogicMP Algorithm
$\begin{split}\mathbf{Q}_{\mathtt{S}}(1)&=\exp(\Phi_{\mathtt{S}}(1)+w_{1}e_{1}+w_{2}e_{3})/\mathbf{Z}_{\mathtt{S}}\,,\\ \mathbf{Q}_{\mathtt{S}}(0)&=\exp(\Phi_{\mathtt{S}}(0)+w_{1}e_{2}+w_{2}e_{4})/\mathbf{Z}_{\mathtt{S}}\,,\\ \end{split}$
@@
$\mathtt{S}(x)\wedge\mathtt{F}(x,y)\implies\mathtt{S}(y)$ $e_{1}=$einsum(“$x,xy\to y$”, $\mathbf{Q}_{\mathtt{S}}(1)$, $\mathbf{Q}_{\mathtt{F}}(1)$) EN of the true ground premises $\mathtt{S}(x)\wedge\mathtt{F}(x,y)$
$\neg\mathtt{S}(y)\wedge\mathtt{F}(x,y)\implies\neg\mathtt{S}(x)$ $e_{2}=$einsum(“$y,xy\to x$”, $1-\mathbf{Q}_{\mathtt{S}}(1)$, $\mathbf{Q}_{\mathtt{F}}(1)$) EN of the true ground premises $\neg\mathtt{S}(y)\wedge\mathtt{F}(x,y)$
$\mathtt{C}(x)\implies\mathtt{S}(x)$ $e_{3}=$einsum(“$x\to x$”, $\mathbf{Q}_{\mathtt{C}}(1)$) EN of the true ground premises $\mathtt{C}(x)$
$\neg\mathtt{C}(x)\implies\neg\mathtt{S}(x)$ $e_{4}=$einsum(“$x\to x$", $1-\mathbf{Q}_{\mathtt{C}}(1)$) EN of the true ground premises $\neg\mathtt{C}(x)$
@@
Figure 15: Facts (top) and predictions (bottom) on the Smoke dataset ($\blacksquare$/$\blacksquare$ for the true/false facts).

### v4

@@ 3 Efficient Mean-field Iteration via LogicMP
$\phi_{f}(v_{g})$ $v_{\mathtt{C}(e_{1},e_{3})}$
useful
0 11
$g_{-\mathtt{C}(e_{1},e_{3})}$
$(0,0)$ $1=1$
@@ 3.2 Parallel Aggregation using Einstein Summation
[…]
@@ 5.2 Encoding FOLCs over Relational Graphs
MLN
MCMC (Richardson & Domingos, 2006) .53 - - - - - - - - - - - - - - - - -
@@ 5.3 Encoding FOLCs over Text
Our Method. For a fair comparison, we also use BLSTM as the backbone and stack LogicMP on BLSTM to integrate the following FOLCs used in LogicDist. (1) adjacent rule: The BIOES schema contains constraints for adjacent labels, e.g., the successive label of B-PER cannot be O-PER. We explicitly declare the constraints as several adjacent logic rules, such as $\forall i:\mathtt{label}(i)\in\{\text{B/I-PER}\}\Leftrightarrow\mathtt{label}(i+1)\in\{\text{I/E-PER}\}$, where $\mathtt{label}(i)$ is the multi-class predicate of $i$-th token label (see the extension for multi-class predicate in App. E). (2) list rule: we exploit a task-specific rule to inject prior knowledge from experts. Specifically, named entities in a list are likely to be in the same categories, e.g., “Barcelona” and “Juventus” in “1. Juventus, 2. Barcelona, 3. …”. The corresponding FOLC is $\forall i,j:\mathtt{label}(i)\in\{\text{B/I/E-LOC}\}\wedge\mathtt{samelist}(i,j)\Leftrightarrow\mathtt{label}(j)\in\{\text{B/I/E-LOC}\}$, where $\mathtt{samelist}(i,j)$ indicates the coexistence of two tokens in a list.
@@ Appendix A Mean-field Iteration Equation of MLN Inference
[…]
@@
[…]
@@
[…]
@@ Appendix B Terminology
[…]
@@
[…]
@@ Appendix C Proof of Theorem: Message of Clause Considers True Premise Only
[…]
@@ Appendix D Proof of Theorem: Message of CNF = $\sum$ Message of Clause
[…]
@@ Appendix G Einsum Examples
One may notice that the current implementations of the Einsum function are not available when the target matrix has external arguments that are not in the input matrices, e.g., $\text{``}a\to ab\text{''}$. We tackle this by a post-processing function for the output of Einsum.
@@ Appendix I An Example to Illustrate the LogicMP Algorithm
[…]
@@
$\mathtt{S}(x)\wedge\mathtt{F}(x,y)\implies\mathtt{S}(y)$ $e_{1}=$einsum($x,xy\to y$”, $\mathbf{Q}_{\mathtt{S}}(1)$, $\mathbf{Q}_{\mathtt{F}}(1)$) EN of the true ground premises $\mathtt{S}(x)\wedge\mathtt{F}(x,y)$
$\neg\mathtt{S}(y)\wedge\mathtt{F}(x,y)\implies\neg\mathtt{S}(x)$ $e_{2}=$einsum($y,xy\to x$”, $1-\mathbf{Q}_{\mathtt{S}}(1)$, $\mathbf{Q}_{\mathtt{F}}(1)$) EN of the true ground premises $\neg\mathtt{S}(y)\wedge\mathtt{F}(x,y)$
$\mathtt{C}(x)\implies\mathtt{S}(x)$ $e_{3}=$einsum($x\to x$”, $\mathbf{Q}_{\mathtt{C}}(1)$) EN of the true ground premises $\mathtt{C}(x)$
$\neg\mathtt{C}(x)\implies\neg\mathtt{S}(x)$ $e_{4}=$einsum($x\to x$", $1-\mathbf{Q}_{\mathtt{C}}(1)$) EN of the true ground premises $\neg\mathtt{C}(x)$
@@
Figure 15: Facts (top) and predictions (bottom) on the Smoke dataset ($\blacksquare$$\blacksquare$ for the true/false facts).